# SPIRAL: Self-Play on Zero-Sum Games Incentivizes Reasoning via Multi-Agent Multi-Turn Reinforcement Learning

### v2

@@ C.2 Comprehensive Benchmark Results
Qwen3-4B-Base 65.8 10.0 3.3 33.3 50.0 24.3 30.6 47.2 33.1
+ SFT 76.0 16.7 13.3 38.4 50.0 31.2 33.0 48.8 38.4
+ Mistral Opponent (KuhnPoker) 64.0 6.7 6.7 29.8 50.0 26.1 35.6 43.6 32.8
+ Gemini Opponent (KuhnPoker) 69.2 10.0 13.3 33.8 50.0 33.8 35.3 55.5 37.6
@@
+ SPIRAL (KuhnPoker) 76.4 13.3 10.0 38.4 57.5 42.4 37.0 56.5 41.4
+ SPIRAL (Negotiation) 76.4 10.0 10.0 41.2 57.5 36.0 34.7 54.2 40.0
+ SPIRAL (TicTacToe+KuhnPoker) 76.2 13.3 16.7 40.7 60.0 41.5 35.7 57.2 42.7
+ SPIRAL (TicTacToe+KuhnPoker+Negotiation) 74.2 20.0 16.7 39.4 55.0 38.6 36.9 57.7 42.3
@@
DeepSeek-Distill-Qwen-7B 90.8 46.7 36.7 56.9 92.5 48.2 48.6 57.1 59.7
+ SPIRAL (Multi-Game) 93.0 43.3 46.7 57.9 92.5 51.1 49.6 58.9 61.7
@@
Appendix D Case Study Methodology
@@
D.1 Data Collection Framework
@@
D.2 Bottom-Up Pattern Discovery Process
@@
D.3 Cross-Domain Transfer Quantification
@@
D.4 Pattern Evolution Analysis
@@
D.5 Concrete Transfer Example Identification
@@
D.6 Pattern Classification at Scale
@@ D.7 Validation Methodology
Appendix E Game Environment Specifications
@@
E.1 Training Game Environments
@@
Simple Negotiation develops multi-constraint optimization through resource trading. Two players exchange Wood and Gold tokens, with opposing utility functions creating natural tension. Each player aims to maximize portfolio value through proposals and counteroffers. Success requires understanding opponent preferences, planning multi-step trades, and strategic communication. We expect these skills to enhance optimization problems and multi-constraint reasoning tasks.
E.2 Out-of-Distribution Evaluation Games
@@
Connect Four adds directional constraints to spatial reasoning. Players drop discs into columns with gravity, aiming to connect four vertically, horizontally, or diagonally. The vertical placement mechanic tests whether TicTacToe strategies generalize to games with restricted movement options.
@@
Liar’s Dice combines probability with deception detection. Players make escalating claims about hidden dice totals and challenge suspected bluffs. Success requires probability calculation, opponent modeling, and strategic deception, testing whether Kuhn Poker skills scale to multi-round bluffing games.

### v3

@@ E.4 Comprehensive Benchmark Results
Qwen3-4B-Base 73.4 9.6 6.2 33.3 42.4 29.4 30.6 47.2 34.0
+ SFT (Kuhn) 74.0 11.0 10.4 36.7 48.6 36.8 33.0 48.8 37.4
+ SFT (Multi, 25k) 74.2 13.7 11.7 37.6 51.1 40.1 37.8 51.3 39.7
+ SFT (Multi, 52k) 73.4 16.2 11.9 39.9 48.7 38.7 38.2 51.2 39.7
+ Mistral Opponent (KuhnPoker) 64.0 4.3 2.1 29.8 31.6 26.1 35.6 43.6 29.6
+ Gemini Opponent (KuhnPoker) 69.2 5.2 4.7 33.8 29.8 33.8 35.3 55.5 33.4
@@
+ SPIRAL (KuhnPoker) 76.4 18.2 15.6 38.4 61.2 42.4 37.0 57.7 43.4
+ SPIRAL (Negotiation) 75.6 11.7 10.2 38.1 51.7 39.3 36.7 57.0 40.0
+ SPIRAL (TicTacToe+KuhnPoker) 76.2 11.4 10.7 40.7 57.2 41.5 35.7 57.2 41.3
+ SPIRAL (Multi-Game) 78.2 19.7 13.3 41.8 61.6 42.6 40.1 58.5 44.5
Qwen3-8B-Base Family
Qwen3-8B-Base 77.0 12.1 11.2 33.5 50.6 38.2 38.0 55.7 39.5
+ SFT (Multi, 25k) 82.8 19.9 15.6 45.9 63.5 40.8 41.6 58.8 46.1
+ SFT (Multi, 52k) 81.8 22.6 14.0 48.8 61.6 40.0 43.1 57.4 46.2
+ SPIRAL (Multi-Game) 86.6 26.2 16.8 49.6 65.2 46.3 44.6 61.1 49.6
Octothinker-8B-Base Family
Octothinker-8B-Base 65.6 1.7 0.5 26.6 33.5 25.7 22.1 30.8 25.8
+ SFT (Multi, 25k) 66.0 3.3 3.8 23.9 31.0 23.8 24.9 39.1 27.0
+ SFT (Multi, 52k) 66.4 4.2 3.6 21.9 31.6 24.6 24.9 40.8 27.3
+ SPIRAL (Multi-Game) 68.6 5.3 4.8 33.7 43.2 32.0 33.8 49.3 33.8
Llama-3.1-8B-Instruct Family
Llama-3.1-8B-Instruct 46.4 4.6 0.7 13.8 23.3 22.8 30.2 49.1 23.9
+ SFT (Multi, 25k) 51.8 4.6 0.7 19.1 23.3 21.7 30.0 48.9 25.0
+ SFT (Multi, 52k) 51.0 4.3 0.0 21.5 23.2 21.1 30.2 49.1 25.1
+ SPIRAL (Multi-Game) 49.8 4.9 1.8 17.3 26.0 24.6 32.2 50.4 25.9
@@
DeepSeek-Distill-Qwen-7B 90.8 53.0 39.5 56.9 89.3 48.2 48.6 57.1 60.4
+ SFT (Multi) 91.8 49.3 36.6 52.4 88.2 48.2 44.5 55.6 58.3
+ SPIRAL (Multi-Game) 93.0 54.1 40.8 57.9 89.3 51.1 49.6 58.9 61.8
@@
We curated an additional 27k dataset with Qwen3-32B self-play and evaluated SFT on a total of 52k trajectories but training with 1 epoch. As shown in the table, doubling the SFT data yields no major improvement (e.g., Qwen3-4B Average 39.7% vs 39.7%), while SPIRAL consistently outperforms both SFT baselines. This confirms that the benefits of SPIRAL stem from the reinforcement learning dynamic rather than simply dataset size.
E.5 Statistical Robustness
To address concerns regarding statistical significance, we re-ran our main experiments with 3 random seeds (seeds 14, 42, 100). As shown in Table˜10, SPIRAL consistently outperforms the SFT baseline with narrow confidence intervals, confirming the robustness of our gains.
Table 10: Comprehensive performance comparison across multiple seeds. SPIRAL demonstrates consistent improvement over baselines across diverse benchmarks.
Model Math500 AIME24 AIME25 Olympiad AMC-23 Minerva GPQA-D MMLU-Pro Average
Qwen3-4B-Base 73.4 9.6 6.2 33.3 42.4 29.4 30.6 47.2 34.0
+ SFT-Multi $74.0\pm 0.6$ $2.4\pm 1.5 11$$1.2\pm 1.1$ $37.8\pm 0.3$ $52.2\pm 0.8$ $40.7\pm 0.5$ $37.7\pm 1.2$ $50.9\pm 0.5$ $39.6\pm 0.4$
+ SPIRAL-Multi (Ours) $\mathbf{78.7\pm 2.0}$ $\mathbf{18.8\pm 2.5}$ $\mathbf{15.0\pm 1.3}$ $\mathbf{41.8\pm 1.3}$ $\mathbf{62.0\pm 1.6}$ $\mathbf{42.1\pm 1.3}$ $\mathbf{39.1\pm 3.1}$ $\mathbf{58.4\pm 0.5}$ $\mathbf{44.5\pm 0.5}$
E.6 Game Trajectory Statistics
In Table˜11, we added a comprehensive table with the average number of game lengths, reasoning tokens per step, P1/P2 self-play win rates, and average win-rate and win-rate per game against Gemini-2.0-Flash.
Table 11: Game trajectory statistics across training checkpoints. We observe increasing game length and reasoning tokens alongside improved win rates against the fixed opponent.
Training Avg. Game Round Avg. Reasoning Self-Play Self-Play Avg. Win-Rate Win-Rate vs. Gemini Win-Rate vs. Gemini Win-Rate vs. Gemini
Steps (Moves) Tokens / Step P1 Win-Rate P2 Win-Rate vs. Gemini (TicTacToe) (Kuhn Poker) (Simple Negotiation)
Step 0 1.69 4061 42.7% 57.3% 12.5% 12.5% 6.25% 18.8%
Step 128 7.63 1609 48.7% 51.3% 28.5% 16.7% 37.5% 31.2%
Step 256 8.47 1755 53.2% 46.8% 51.8% 42.9% 68.8% 43.8%
Step 384 9.43 1921 59.1% 40.9% 66.7% 75.0% 68.8% 56.3%
Step 400 9.55 2032 62.4% 37.6% 67.4% 83.3% 62.5% 56.3%
E.7 SPIRAL as part of the mid-training stage
[…]
Table 12: Performance comparison of SPIRAL integrated into different training stages. SPIRAL acts as a robust performance booster both before and after RLVR.
Model Math500 AIME24 AIME25 Olympiad AMC-23 Minerva GPQA-D MMLU-Pro Average
Qwen3-4B-Base 73.4 9.6 6.2 33.3 42.4 29.4 30.6 47.2 34.0
RLVR (Math) 83.0 18.4 15.6 44.6 62.8 43.4 43.7 56.8 46.0
SPIRAL 78.2 19.7 13.3 41.8 61.6 42.6 40.1 58.5 44.5
SPIRAL $\to$ RLVR 84.2 23.1 17.2 45.2 59.6 42.1 43.2 57.4 46.5
RLVR $\to$ SPIRAL 86.1 22.6 18.1 46.0 62.5 44.3 44.7 60.8 47.9
Our results in Table˜12 validate SPIRAL’s effectiveness on integrating as part of the mid-training: SPIRAL $\to$ RLVR outperforms the Base $\to$ RLVR baseline, while applying SPIRAL as a post-RLVR stage (RLVR $\to$ SPIRAL) yields the highest performance (Avg 48.1). Furthermore, on the stronger DeepSeek-Distill-Qwen-7B, SPIRAL-Multi improves the average score to 59.3, whereas standard SFT leads to performance regression. These findings demonstrate that SPIRAL serves as a robust booster that mitigates the alignment tax often seen in standard fine-tuning stages.
E.8 Instruct Model Results
To better study our method’s effectiveness, we perform additional experiments with Qwen3-4B-Instruct-2507.
As shown in the Table˜13, standard SFT on reasoning trajectories slightly degrades performance. In contrast, SPIRAL successfully reverses this trend. SPIRAL achieves a 2% improvement over the original model and a substantial average gain of 3.6% over the SFT baseline, reaching a total average score of 75.51%. This demonstrates that self-play can refine genuine reasoning capabilities even in models that have strong performance.
Table 13: Performance comparison on Qwen3-4B-Instruct. While standard SFT leads to regression, SPIRAL improves reasoning capabilities across most benchmarks.
Model Math500 AIME24 AIME25 Olympiad AMC-23 Minerva GPQA-D MMLU-Pro Average
Qwen3-4B-Instruct (Base) 91.2 64.6 47.4 89.6 82.1 87.6 62.0 69.6 74.10
+ SFT-Multi 90.8 59.3 46.1 86.4 79.6 82.3 61.4 69.1 71.88
+ SPIRAL-Multi (Ours) 93.5 67.9 49.8 88.8 83.3 85.5 61.8 71.5 75.91
Appendix F Case Study Methodology
@@
F.1 Data Collection Framework
@@
F.2 Bottom-Up Pattern Discovery Process
@@
F.3 Cross-Domain Transfer Quantification
@@
F.4 Pattern Evolution Analysis
@@
F.5 Concrete Transfer Example Identification
@@
F.6 Pattern Classification at Scale
@@ F.7 Validation Methodology
Appendix G Game Environment Specifications
@@
G.1 Training Game Environments
@@
Simple Negotiation is a game that develops multi-constraint optimization skills through resource trading. Two players exchange Wheat, Wood, Sheep, Brick, and Gold tokens. Since the utility of these resources varies for each player, there is a natural incentive to trade. Each player aims to maximize their portfolio’s value by making proposals and counteroffers. Success requires understanding an opponent’s preferences, planning multi-step trades, and communicating strategically. We expect these skills to improve performance on optimization problems and multi-constraint reasoning tasks.
G.2 Out-of-Distribution Evaluation Games